1. Assume that a certain simplified economy produces only two goods, X and Y, with given resources and technology. The following table gives the various possible combinations of the production of the two goods (all units are measured in millions of tons).Opportunity Cost ofProduction Possibility Good X Good Y Good X
A 0 100
B 2 90
C 4 60
D 6 20
a) Calculate the opportunity cost of the production of good X at each point. What law does the trend in those values exhibit?
b) What changes are required for this economy to shift the PPF outward?

1
Expert's answer
2021-04-12T13:04:35-0400


At point B the opportunity cost will be;

"=\frac{90-100}{2-0}=\frac{-10}{2}=-5"


At point C the opportunity cost will be;

"=\frac{60-90}{4-2}=\frac{-30}{2}=-15"


At point D the opportunity cost will be;

"=\frac{20-60}{6-4}=\frac{-40}{2}=-20"

Considered from point D towards A, the values exhibit the Law of increasing opportunity cost. However, considered from point A towards D, the values show a decreasing tendency and thus do not obey the law of increasing opportunity cost.


(b)In order to shift the PPF outward, the following can be done:

  • Increasing total amount of available production factors like labor and capital.
  • Advancements in technology.If new technologies are developed that enable goods to be produced with fewer production factors, the economy’s capital will essentially experience a rise in purchasing power parity. Thus, the economy will be able to produce more at any point along the frontier, meaning that the frontier has effectively shifted outwards